# Threshold Graphs are Globally Synchronizing

### v4

@@ 5.2. Phasor Geometry of Closed Twins
Open twins, defined by $N(i)=N(j)$, will not be needed in this work.
@@
Let $\boldsymbol{\theta}$ be a second-order stationary point of the energy function $E_{G}(\boldsymbol{\theta})$. If nodes $a$ and $b$ form a pair of closed twins, then $\theta_{a}=\theta_{b}$.
@@
By Corollary 5.2 and the fact that $\mu_{a},\mu_{b}\geq 0$, we conclude that $\boldsymbol{v}_{a}=\boldsymbol{v}_{b}$, and therefore $\theta_{a}=\theta_{b}$. ∎
Clearly, this is a case where geometric twins defined in (5.3) formed. Moreover, Corollary 5.5 implies immediately that windmill graphs are globally synchronizing.
@@
Let $\boldsymbol{\theta}$ be a second-order stationary point of $E_{G}(\boldsymbol{\theta})$, and let $G$ be a windmill graph. Within each clique of $G$, all non-central nodes have identical closed neighborhoods and therefore form closed twins. By Corollary 5.5, all nodes in each clique must share the same phase. Contracting each clique into a single node yields a star graph, which is globally synchronizing. Hence all nodes of $G$ have the same phase. ∎
@@ 5.3. Synchronous Pendant Extension
In this section, we present a scenario in which a group of nodes satisfy the phasor relations of Lemma 5.1 without being closed twins. More explicitly, the main result of this section is Lemma 5.8, which shows that even when a set of nodes are not twins, they may still be forced to synchronize at any second-order stationary point, provided that their distinct neighbors form a synchronous pendant whose combined phasor contribution collapses to a single, non-adversarial direction.
The next lemma, Lemma 5.7, serves as a preparatory step toward establishing Lemma 5.8.
@@
Let $G=(V,E)$ be a graph, and let $W\subseteq V$ be partitioned into
@@
subject to the following structural conditions:
  1. (i)
    Every node in $Q$ has all its neighbors inside $S$, i.e., $N(i)\subseteq S$ for all $i\in Q$;
  2. (ii)
    $S$ induces a clique in $G$;
  3. (iii)
    For each $i\in S$, its neighborhood is $P\cup N_{Q}(i)\cup(S\setminus\{i\})$, where $N_{Q}(i)=\{\,q\in Q:q\sim i\,\}$.
Assume further that the configuration $(\boldsymbol{v}_{i})_{i\in V}$ is a second-order stationary point of the Kuramoto energy (4), and that $\boldsymbol{v}_{i}=\boldsymbol{v}$ for all $i\in Q\uplus S$. Then the phasor sum over $P$ satisfies
@@
$N(i)=Q\uplus(S\setminus\{i\})\uplus P$ (15)
@@
$\sum_{j\in Q\uplus(S\setminus\{i\})\uplus P}\boldsymbol{v}_{j}=\mu_{i}\boldsymbol{v}_{i}.$
@@
$\bigl(|Q|+|S|-1\bigr)\,\boldsymbol{v}_{i}+\sum_{j\in P}\boldsymbol{v}_{j}=\mu_{i}\boldsymbol{v}_{i},\qquad\mu_{i}\geq 0.$
@@
$\sum_{j\in P}\boldsymbol{v}_{j}=\bigl(\mu_{i}-(|Q|+|S|-1)\bigr)\boldsymbol{v}_{i},\qquad\mu_{i}\geq 0.$
We now show that $\mu_{i}-(|Q|+|S|-1)<0$ is impossible, because in this case there would exist a direction along which the energy decreases in a neighborhood of $\boldsymbol{\theta}$.
To see this, let $\boldsymbol{v}$ be the $n$-dimensional indicator vector of $Q\uplus S$, i.e., $v_{i}=1$ for $i\in Q\uplus S$ and $v_{i}=0$ otherwise. Then
$\boldsymbol{v}^{T}H\boldsymbol{v}$ $=\sum_{i\in Q\uplus S}\sum_{j\in P}\cos(\theta_{i}-\theta_{j})$ (17)
@@
where the second equality uses the fact that there are no edges between $Q$ and $P$. Clearly, if $\mu_{i}-(|Q|+|S|-1)<0$, then the above quantity is negative, which contradicts the second-order stationarity of $\boldsymbol{\theta}$. ∎
Figure 11. Illustration of Lemma 5.7. The green vertices correspond to nodes for which $\mu_{i}\boldsymbol{v}_{i}=\sum_{j\in N(i)}\boldsymbol{v}_{j}$ holds with $\mu_{i}\geq 0$. The large circular node represents the set $P$ of vertices. Gray edges indicate synchronization interactions between the corresponding nodes.
@@
Let $G=(V,E)$ be a graph, and suppose there exists a subset of nodes $W\subseteq V$ partitioned into four disjoint subsets:
$W=Q\uplus S_{1}\uplus S_{2}\uplus P.$
under the following structural assumptions:
(i) For all $i\in Q$, we have $N(i)\subseteq S_{1}$;
(ii) The subgraph $G[S_{1}\uplus S_{2}]$ is a complete graph;
(iii) For any $i\in S_{1}\uplus S_{2}$, its neighbourhood can be decomposed into three parts: the set $P$, its neighbours inside $Q$, and the rest of the clique $S_{1}\uplus S_{2}$. Formally, if we denote $Q_{i}=N(i)\cap Q$, then
$N(i)\;=\;P\uplus Q_{i}\uplus\bigl((S_{1}\uplus S_{2})\setminus\{i\}\bigr).$
Equivalently,
$N(i)\setminus\bigl(Q\uplus(S_{1}\uplus S_{2})\bigr)\;=\;P.$
Assume further that all nodes in $Q\uplus S_{1}$ are synchronized, i.e.,
$\boldsymbol{v}_{i}=\boldsymbol{v}_{j},\quad\forall i,j\in Q\uplus S_{1}.$
Then the set $S_{1}\uplus S_{2}$ synchronize.
@@
Observe that the set $S_{1}$ induces a fully-connected subgraph, and $Q$ is a synchronous pendant to $S_{1}$, then according to Lemma 5.8, we have
$\forall i\in S_{1},\quad\boldsymbol{v}_{i}\Uparrow\sum_{k\in S_{2}\uplus P}\boldsymbol{v}_{k}\quad\text{or}\quad\sum_{k\in S_{2}\uplus P}\boldsymbol{v}_{k}=\boldsymbol{0}.$
Combine this with the fact
$\boldsymbol{v}_{i}=\boldsymbol{v}_{j}\quad\forall i\in S_{1},j\in S_{2},$
we obtain
$\forall i\in S_{1},\quad\boldsymbol{v}_{i}\Uparrow\sum_{k\in S_{2}\uplus P}\boldsymbol{v}_{j}+\sum_{k\in S_{1}\setminus\{i\}}\boldsymbol{v}_{j}\quad\text{or}\quad\sum_{k\in S_{2}\uplus P}\boldsymbol{v}_{k}+\sum_{k\in S_{1}\setminus\{i\}}\boldsymbol{v}_{k}=\boldsymbol{0}.$
Equivalently, we can write
$\forall i\in S_{1},\quad\boldsymbol{v}_{i}\Uparrow\sum_{k\in S_{1}\uplus P\uplus S_{2}\setminus\{i\}}\boldsymbol{v}_{k}\quad\text{or}\quad\sum_{k\in S_{1}\uplus P\uplus S_{2}\setminus\{i\}}\boldsymbol{v}_{k}=\boldsymbol{0}.$ (18)
Another observation is on nodes in $S_{2}$ that
$\forall i\in S_{2},\quad\boldsymbol{v}_{i}\Uparrow\sum_{k\in S_{2}\setminus\{i\}}\boldsymbol{v}_{k}+\sum_{j\in S_{1}\uplus P}\boldsymbol{v}_{k}\quad\text{or}\quad\sum_{k\in S_{2}\setminus\{i\}}\boldsymbol{v}_{k}+\sum_{j\in S_{1}\uplus P}\boldsymbol{v}_{k}=\boldsymbol{0}$
since every node in $S_{2}$ attains stable equilibrium in $\boldsymbol{\theta}$. Equivalently, we can write
@@
Combined the facts (18) and (19), we get nodes in $S_{1}\uplus S_{2}$ are geometric stable closed twins. ∎
Clearly, this is another situation in which geometric twins, as defined in Definition 5.3, appear in addition to closed twins at the second-order stationary point.
Figure 12. Illustration of Lemma 5.8, illustrating how the synchronized block $Q\uplus S_{1}$ extends to include $S_{2}$ through propagation enabled by the specific local structure. The green vertices correspond to nodes for which $\mu_{i}\boldsymbol{v}_{i}=\sum_{j\in N(i)}\boldsymbol{v}_{j}$ holds with $\mu_{i}\geq 0$. The gray-shaded nodes are synchronized; an edge with a gray background indicates synchronization between the corresponding nodes. Each orange arc separates a small green node from a group of nodes it connects to, indicating that the vector of the green node is in the same direction as the sum of the vectors of the nodes on the opposite side of the arc.
5.4. Twin Attachment
If a set of nodes all have identical neighborhoods, they form a group of closed twins. When an additional node is connected exclusively to this group, the entire collection of nodes, must take the same phase at any second-order stationary point.
Lemma 5.9.
Let $\boldsymbol{\theta}$ be a second-order stationary point of $E_{G}(\boldsymbol{\theta})$. If $S\subseteq V$ is a set of closed twins in $G$ and $A\in V$ satisfies $N(A)=S$, then $\theta_{i}=\theta_{j}$ for all $i,j\in S\cup\{A\}$.
Proof.
Since $\boldsymbol{\theta}$ is a second-order stationary point of $E_{G}(\boldsymbol{\theta})$ and $S$ forms a set of closed twins, Lemma 5.5 yields
$\theta_{i}=\theta_{j}\qquad\text{for all }i,j\in S.$
Because $N(A)=S$, Lemma 4.4 implies
$\theta_{A}=\theta_{i}\qquad\text{for all }i\in S.$
Hence all nodes in $S\cup\{A\}$ synchronize. ∎
Figure 13. Illustration of Lemma 5.9. All green nodes attain stable equilibrium. The set $S$ (middle row) forms a set of false twins, and node AA (top) has neighborhood exactly $S$. Thin lines represent the edges in the original graph. The conclusion is that all green nodes must synchronize, indicated by the thick edges in the right diagram.
@@ 6.1. A Prototype Example
Let us take a look at the threshold graph corresponding to the sequence 01010101. The first vertex is denoted by $A$, and we label the vertices $A$ through $I$ in the order in which they appear in the sequence, with each bit determining whether the added vertex is isolated (0) or dominating (1).
Let $G$ be the threshold graph corresponding to the sequence 01010101. To prove that $G$ is globally synchronizing, we show that every second-order stationary point $\boldsymbol{\theta}$ of $E_{G}(\boldsymbol{\theta})$ is synchronous; that is, $\theta_{1}=\cdots=\theta_{n}$. The following figure illustrates how synchronization must propagate, step by step, in any second-order stationary configuration on this graph.
@@
Figure 14. Synchronization propagates in eight steps, forcing every second-order stationary point to be synchronous.
Step (1) We begin with the edge between $H$ and $I$. Applying Lemma 4.4 to node $H$ shows that, since $\boldsymbol{\theta}$ is a second-order stationary point, we must have $\boldsymbol{v}_{H}\Uparrow\boldsymbol{v}_{I}$. Equivalently, nodes $H$ and $I$ synchronize.
Step (2) Observe that $N(G)=N(I)\cup H$ and $\theta_{H}=\theta_{I}$, applying Lemma by letting $Q=\{H\}$, $S_{1}=\{I\}$, $S_{2}=\{G\}$, $P=\{F,D,B,A,C,E\}$, we obtain $\theta_{G}=\theta_{I}$. Intuitively, the only difference between the neighborhoods of $G$ and $H$ is that $H$ has one additional neighbor, namely $I$. This extra neighbor does not obstruct the synchronization between $G$ and $I$, because $H$ is already synchronized with $I$ and therefore behaves as a synchronous pendant at this stage.
Step (3) Note that node $F$ connects only to $G$ and $I$, which are synchronized as established in Step (2). Therefore, by Lemma 5.9, we have $\boldsymbol{v}_{F}=\boldsymbol{v}_{G}=\boldsymbol{v}_{I}$.
Step (4) Observe that nodes $F$ and $H$ connect only to $G$ and $I$, and the four-node set $\{F,H,G,I\}$ behaves as a synchronizing group. Furthermore, ignoring the leaf-like nodes $F$ and $H$ and edges in the induced clique on $\{E,G,I\}$, the two nodes $G$ and $I$ share exactly the same neighbors:
$N(G)\setminus(\{F\}\cup\{E,I\})\;=\;N(I)\setminus(\{F,H\}\cup\{E,G\})\;=\;N(E)\setminus\{G,I\}\;=\;\{D,B,A,C\}.$
Therefore, by applying Lemma 5.8 with $Q=\{F,H\}$, $S_{1}=\{G,I\}$, $S_{2}=\{E\}$, and $P=\{D,B,A,C\}$, we conclude that the nodes $E$, $G$, and $I$ form a synchronizing group.
Step (5) Note that node $D$ connects to $E$, $G$, and $I$, which are synchronized as established in Step (4). Therefore, by Lemma 5.9, we have $\boldsymbol{v}_{D}=\boldsymbol{v}_{E}=\boldsymbol{v}_{G}=\boldsymbol{v}_{I}$.
Step (6) Observe that nodes $D$, $F$, and $H$ connect only to $E$, $G$, and $I$, and therefore the six-node set $\{D,F,H,E,G,I\}$ behaves as a synchronizing group. Furthermore, ignoring the leaf-like nodes $D$, $F$, and $H$ and edges in the induced clique on $\{C,E,G,I\}$, the four nodes $C$, $E$, $G$, and $I$ share exactly the same neighbors, namely $\{A,B\}$. Therefore, by applying Lemma 5.8 with $Q=\{D,F,H\}$, $S_{1}=\{E,G,I\}$, $S_{2}=\{C\}$, and $P=\{A,B\}$, we conclude that the nodes $C$, $E$, $G$, and $I$ form a synchronizing group.
Step (7) Note that node $B$ connects to $C$, $E$, $G$, and $I$, which are synchronized as established in Step (6). Therefore, by Lemma 5.9, we have $\boldsymbol{v}_{B}=\boldsymbol{v}_{C}=\boldsymbol{v}_{E}=\boldsymbol{v}_{G}=\boldsymbol{v}_{I}$.
Step (8) As in Step (7), Lemma 5.9 implies that $A$ synchronizes with $C$, $E$, $G$, and $I$, since these nodes form a synchronizing group and constitute all neighbors of $A$.
In conclusion, $\boldsymbol{\theta}$ must assign the same angle to every node. Equivalently, every second-order stationary point is synchronous, and hence the threshold graph is globally synchronizing.
@@
Any connected threshold graph can be described by a bit sequence of $1$’s and 0’s, as introduced in Definition 2.1. Since the graph is connected, the last bit must be a $1$. We decompose the sequence into maximal contiguous blocks of $1$’s and 0’s. Let $U_{1},U_{2},\dots,U_{k}$ denote the blocks of $1$’s (indexed from bottom to top), and let $I_{i}$ denote the 0-block that lies immediately above $U_{i}$, if such a block exists. In this notation, the sequence has one of the two forms
@@
depending on whether or not it begins with a block of 0’s. We focus on the first case; the second follows by an entirely analogous argument.
Given that $\boldsymbol{\theta}$ is a second-order stationary point of $E(\boldsymbol{\theta})$, we aim to show that $\theta_{1}=\cdots=\theta_{n}$, and we proceed by induction. We proceed by induction. For the base case, observe that all nodes within a block of $1$’s form a set of closed twins. Hence, at any second-order stationary point $\boldsymbol{\theta}$, the nodes in the last block $U_{k}$ must synchronize, as follows directly from Lemma 5.1.
Assume as the inductive hypothesis that the nodes in the last $m$ blocks of $1$’s, $U_{k-m+1},\dots,U_{k}$, are synchronized. Lemma 5.9 then implies that the adjacent 0-blocks $I_{k-m+1},\dots,I_{k}$ synchronize with them as well. We next show that the preceding $1$-block $U_{k-m}$ also synchronizes with $U_{k-m+1},\dots,U_{k}$. Indeed, define

### v5

@@ 5.2. Phasor Geometry of Closed Twins
Open twins, corresponding to $N(i)=N(j)$, will not be needed in this work.
@@
Let $\boldsymbol{\theta}\in\mathbb{R}^{n}$ be a second-order stationary point of the energy function $E_{G}(\boldsymbol{\theta})$. If nodes $a$ and $b$ form a pair of closed twins (satisfying (10)), then $\boldsymbol{v}_{a}=\boldsymbol{v}_{b}$.
@@
By Corollary 5.2 and the fact that $\mu_{a},\mu_{b}\geq 0$, we conclude that $\boldsymbol{v}_{a}=\boldsymbol{v}_{b}$. ∎
Clearly, this is a case where geometric twins defined in (5.4) formed.
Moreover, Corollary 5.6 implies immediately that windmill graphs are globally synchronizing.
@@
Let $G$ be a windmill graph $W_{p,m}$ with central vertex $c$. Then there exists a partition of the vertex set
$V\setminus\{c\}\;=\;\biguplus_{\ell=1}^{m}S_{\ell}$
such that each induced subgraph $G[S_{\ell}]$ is a clique, and every vertex in $S_{\ell}$ is adjacent to $c$. For each $\ell$, all vertices in $S_{\ell}$ have identical closed neighborhoods and therefore form a class of closed twins. By Corollary 5.6, all vertices in each $S_{\ell}$ must share the same phase. According to Lemma (4.4), we have for $\forall i\in S_{l}$, that
$\boldsymbol{v}_{i}\Uparrow\boldsymbol{v}_{c}+\sum_{j\in S_{l},j\neq i}\boldsymbol{v}_{j}\qquad\text{or}\qquad\boldsymbol{v}_{c}+\sum_{j\in S_{l},j\neq i}\boldsymbol{v}_{j}=\boldsymbol{0}.$
We distinguish two cases.
Case 1. $\boldsymbol{v}_{c}\Uparrow\boldsymbol{v}_{i}+\sum_{j\in S_{\ell},\,j\neq i}\boldsymbol{v}_{j}$. It is then clear that either $\boldsymbol{v}_{c}=\boldsymbol{v}_{i}$ or $\boldsymbol{v}_{c}=-\boldsymbol{v}_{i}$ holds. If $\boldsymbol{v}_{c}=-\boldsymbol{v}_{i}$, then $\boldsymbol{\theta}$ is not a second-order stationary point. Indeed, let $\boldsymbol{x}\in\mathbb{R}^{n}$ be the vector whose $i$-th component equals $1$ if $i\in S_{l}$ and 0 otherwise. Then
$\boldsymbol{x}^{\top}\boldsymbol{H}\boldsymbol{x}$ $=\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1}^{n}\boldsymbol{A}_{ij}\cos(\theta_{i}-\theta_{j})(x_{i}-x_{j})^{2}$
$=\sum_{i\in S_{l}}\sum_{j\in V\setminus S_{l}}\boldsymbol{A}_{ij}\cos(\theta_{i}-\theta_{j})$
$=\sum_{i\in S_{l}}\cos(\theta_{i}-\theta_{c})<0,$
which contradicts the positive semidefiniteness of $\boldsymbol{H}$.
Case 2. $\boldsymbol{v}_{c}+\sum_{j\in S_{\ell},\,j\neq i}\boldsymbol{v}_{j}=\boldsymbol{0}$. Then necessarily $|S_{l}|=2$. In this case, the phasors of the two nodes in $S_{l}$ are opposite to $\boldsymbol{v}_{c}$. By an argument analogous to the one above, the matrix $\boldsymbol{H}$ is not positive semidefinite. ∎
@@ 5.3. Synchronous Pendant Extension
In this section, we present Lemma 5.9 which describes a scenario where a group of nodes forms geometrically stable closed twins 5.4 without being closed twins in the graph-theoretic sense. Before that, we establish Lemma 5.8, which serves as a preparatory step.
@@
Let $G=(V,E)$ be a graph, and let $W\subseteq V$ admit a partition
@@
with the following structural conditions:
(1) Every node in $Q$ has all its neighbors inside $S$, i.e., $N(i)\subseteq S$ for all $i\in Q$;
(2) The vertex set $S\subseteq V(G)$ induces a clique in $G$;
(3) For each $i\in S$, its neighborhood is $P\uplus N_{Q}(i)\uplus(S\setminus\{i\})$, where
$N_{Q}(i)=\{\,q\in Q:q\sim i\,\}.$
Assume further that $\{\boldsymbol{v}_{i}\}_{i\in V}$ is the vector configuration corresponding to any second-order stationary point $\boldsymbol{\theta}$ of the Kuramoto energy (2), and that the induced subgraph on $Q\uplus S$ is synchronized at $\boldsymbol{\theta}$, namely $\boldsymbol{v}_{i}=\boldsymbol{v}$ for all $i\in Q\uplus S$, for some common phasor $\boldsymbol{v}$. Then the phasor sum over $P$ satisfies
@@
$N(i)=N_{Q}(i)\uplus(S\setminus\{i\})\uplus P$ (11)
@@
$\sum_{j\in N_{Q}(i)\uplus(S\setminus\{i\})\uplus P}\boldsymbol{v}_{j}=\mu_{i}\boldsymbol{v}_{i},\qquad\mu_{i}\geq 0$
@@
$\bigl(|N_{Q}(i)|+|S|-1\bigr)\,\boldsymbol{v}_{i}+\sum_{j\in P}\boldsymbol{v}_{j}=\mu_{i}\boldsymbol{v}_{i},\qquad\mu_{i}\geq 0.$ (13)
@@
$\sum_{j\in P}\boldsymbol{v}_{j}=\bigl(\mu_{i}-(|N_{Q}(i)|+|S|-1)\bigr)\boldsymbol{v}_{i},\qquad\mu_{i}\geq 0.$
In fact, $\mu_{i}-(|N_{Q}(i)|+|S|-1)<0$ does not hold, because in this case there exists a direction along which the energy decreases in a neighborhood of $\boldsymbol{\theta}$. To see this, let $\boldsymbol{x}$ be the $n$-dimensional indicator vector of $Q\uplus S$, i.e., $\boldsymbol{x}_{i}=1$ for $i\in Q\uplus S$ and $\boldsymbol{x}_{i}=0$ otherwise. Since $\mu_{i}-(|N_{Q}(i)|+|S|-1)<0$, we have $\langle\sum_{j\in P}\boldsymbol{v}_{j},\boldsymbol{v}_{i}\rangle<0$. Thus
$\boldsymbol{x}^{T}H\boldsymbol{x}$ $=\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1}^{n}\boldsymbol{A}_{ij}\cos(\theta_{i}-\theta_{j})(x_{i}-x_{j})^{2}$ (14)
$=\sum_{i\in Q\uplus S}\sum_{j\in V\setminus(Q\uplus S)}\boldsymbol{A}_{ij}\cos(\theta_{i}-\theta_{j})(x_{i}-x_{j})^{2}$
$=\sum_{i\in S}\sum_{j\in P}\cos(\theta_{i}-\theta_{j})(x_{i}-x_{j})^{2}$
@@
Clearly, if $\mu_{i}-(|N_{Q}(i)|+|S|-1)<0$, then (13) implies that
$\boldsymbol{v}_{i}\cdot\sum_{j\in P}\boldsymbol{v}_{j}<0.$
Consequently, the quantity (14) is negative, which contradicts the second-order stationarity of the state $\boldsymbol{\theta}$. ∎
Figure 10. Illustration of Lemma 5.8. The green vertices correspond to nodes for which $\mu_{i}\boldsymbol{v}_{i}=\sum_{j\in N(i)}\boldsymbol{v}_{j}$ holds with $\mu_{i}\geq 0$. The large circular node represents the set $P$ of vertices. Gray-shaded edges between two nodes indicate that the vectors corresponding to these two nodes are synchronized. The structural relations among the sets $Q$, $S$, and $P$, together with the synchronization of $Q\uplus S$, imply a synchronization in phase between $Q\uplus S$ and the phasor sum over $P$.
@@
Let $G=(V,E)$ be a graph, and let $W\subseteq V$ admit a partition
$W=Q\uplus S_{1}\uplus S_{2}\uplus P$
with the following structural conditions:
(i) Every node in $Q$ has all its neighbors inside $S_{1}$, i.e., $N(i)\subseteq S_{1}$ for all $i\in Q$;
(ii) The vertex set $S_{1}\uplus S_{2}\subseteq V(G)$ induces a clique in $G$;
For each $i\in S_{1}\uplus S_{2}$, its neighborhood is $P\uplus N_{Q}(i)\uplus(S\setminus\{i\})$, where
$N_{Q}(i)=\{\,q\in Q:q\sim i\,\}.$
(iii) For any $i\in S_{1}\uplus S_{2}$, its neighbourhood can be decomposed into three parts: the set $P$, its neighbours inside $Q$, and the rest of the clique $S_{1}\uplus S_{2}$. That is
$N(i)\;=\;P\uplus N_{Q}(i)\uplus\bigl((S_{1}\uplus S_{2})\setminus\{i\}\bigr).$
Assume further that $\{\boldsymbol{v}_{i}\}_{i\in V}$ corresponds to a second-order stationary point $\boldsymbol{\theta}$ of the Kuramoto energy (2), at which the induced subgraph on $Q\uplus S_{1}$ is synchronized, i.e.,
$\boldsymbol{v}_{i}=\boldsymbol{v},\quad\forall i\in Q\uplus S_{1},$
for some common phasor $\boldsymbol{v}$. Then the set $S_{1}\uplus S_{2}$ synchronize.
@@
Step A. Observe that the set $S_{1}$ induces a clique, and $Q\uplus S_{1}$ forms a synchronizing group. Hence by Lemma 5.8, for every $i\in S_{1}$ there exists $\mu_{i}\geq 0$ such that
$\sum_{k\in S_{2}\uplus P}\boldsymbol{v}_{k}=\mu_{i}\,\boldsymbol{v}_{i}.$ (15)
Step B. Since $S_{1}$ is a synchronizing group, we have
$\sum_{k\in S_{1}\setminus\{i\}}\boldsymbol{v}_{k}=(|S_{1}|-1)\boldsymbol{v}_{i}.$
Adding this term to (15) yields
$\sum_{k\in S_{2}\uplus P\uplus(S_{1}\setminus\{i\})}\boldsymbol{v}_{k}=\bigl(\mu_{i}+|S_{1}|-1\bigr)\boldsymbol{v}_{i}.$
where $\mu_{i}+|S_{1}|-1>0$. In other words,
$\forall i\in S_{1},\qquad\boldsymbol{v}_{i}\Uparrow\sum_{k\in S_{2}\uplus P\uplus(S_{1}\setminus\{i\})}\boldsymbol{v}_{k}.$ (16)
Step C. According to Lemma 4.4, nodes in the set $S_{2}$ satisfy
@@
since $\boldsymbol{\theta}$ is a second-order stationary point. Combining (16) and (17), we deduce that all nodes in $S_{1}\uplus S_{2}$ are geometrically stable twins in the sense of Definition 5.4. It then follows from Corollary 5.2 that $S_{1}\uplus S_{2}$ forms a synchronizing group. ∎
Figure 11. Illustration of the proof of Lemma 5.9, illustrating how the synchronized block $Q\uplus S_{1}$ extends to include $S_{2}$ through propagation enabled by the specific local structure. The green vertices correspond to nodes for which $\mu_{i}\boldsymbol{v}_{i}=\sum_{j\in N(i)}\boldsymbol{v}_{j}$ holds with $\mu_{i}\geq 0$. The gray-shaded nodes are synchronized; an edge with a gray background indicates synchronization between the corresponding nodes. Each orange arc separates a small green node from a group of nodes it connects to, indicating that the vector of the green node is in the same direction as the sum of the vectors of the nodes on the opposite side of the arc.
@@ 6.1. A Prototype Example
Let us take a look at the threshold graph corresponding to the sequence 01010101. The first vertex is denoted by $A$, and we label the vertices $A$ through $I$ in the order in which they appear in the sequence.
Let $G$ be the threshold graph corresponding to the sequence 01010101. To prove that $G$ is globally synchronizing, we show that every second-order stationary point $\boldsymbol{\theta}$ of $E_{G}(\boldsymbol{\theta})$ is synchronous; that is, $\theta_{1}=\cdots=\theta_{n}$.
@@
Figure 12. Synchronization propagates in eight steps, forcing every second-order stationary point to be synchronous. Gray-shaded edges represent synchronizing relations that have been established at each step.
Step (1) We begin with the edge between $H$ and $I$. Applying Lemma 4.4 to node $H$ shows that, since $\boldsymbol{\theta}$ is a second-order stationary point, we must have $\boldsymbol{v}_{H}\Uparrow\boldsymbol{v}_{I}$. Equivalently,
$\theta_{H}=\theta_{I}.$
Step (2) Observe that $N(G)=N(I)\cup\{H\}$ and $\theta_{H}=\theta_{I}$, applying Lemma 5.9 by letting $Q=\{H\}$, $S_{1}=\{I\}$, $S_{2}=\{G\}$, $P=\{F,D,B,A,C,E\}$, we obtain
$\theta_{G}=\theta_{I}.$
Intuitively, the only difference between the neighborhoods of $G$ and $I$ is that $I$ has one additional neighbor, namely $H$. This extra neighbor does not obstruct the synchronization between $G$ and $I$, because $H$ is already synchronized with $I$ and therefore behaves as a synchronous pendant at this stage.
Step (3) Note that node $F$ connects only to $G$ and $I$, which are synchronized as established in Step (2). Therefore, by Lemma 4.4, we have
$\theta_{F}=\theta_{G}=\theta_{I}.$
Step (4) Applying Lemma 5.9 with $Q=\{F,H\}$, $S_{1}=\{G,I\}$, $S_{2}=\{E\}$, and $P=\{D,B,A,C\}$, we conclude
$\theta_{E}=\theta_{G}=\theta_{I}.$
Step (5) Note that node $D$ connects to $E$, $G$, and $I$, which are synchronized as established in Step (4). Therefore, by Lemma 4.4, we have
$\theta_{D}=\theta_{E}=\theta_{G}=\theta_{I}.$
Step (6) Applying Lemma 5.9 with $Q=\{D,F,H\}$, $S_{1}=\{E,G,I\}$, $S_{2}=\{C\}$, and $P=\{A,B\}$, we conclude
$\theta_{C}=\theta_{E}=\theta_{G}=\theta_{I}.$
Step (7) Note that node $B$ connects to $C$, $E$, $G$, and $I$, which are synchronized as established in Step (6). Therefore, by Lemma 4.4, we have
$\theta_{B}=\theta_{C}=\theta_{E}=\theta_{G}=\theta_{I}.$
Step (8) As in Step (7), Lemma 4.4 implies that
$\theta_{A}=\theta_{C}=\theta_{E}=\theta_{G}=\theta_{I}$
since these nodes form a synchronizing group and constitute all neighbors of $A$.
In conclusion, every second-order stationary point $\boldsymbol{\theta}$ is a fully synchronized state. Hence the threshold graph corresponding to 01010101 is globally synchronizing.
@@
Any connected threshold graph can be described by a bit sequence of $1$’s and 0’s, where the last bit must be a $1$; otherwise, the threshold graph is not connected. For simplicity, we decompose the sequence into maximal contiguous blocks of $1$’s and 0’s:
@@
where $U_{i}$ denotes a subsequence of $1$’s and $I_{i}$ denotes a subsequence of 0’s. Without loss of generality, we focus on threshold graphs $G=(V,E)$ of the first sequence type; the second case follows by an entirely analogous argument.
Let $\boldsymbol{\theta}$ be a second-order stationary point of the energy function $E_{G}(\boldsymbol{\theta})$ of the homogeneous Kuramoto model on $G$. We prove that $\boldsymbol{\theta}$ is a fully synchronized state by induction on $m$, where $m$ denotes the number of $1$-blocks counted from the end of the threshold sequence.
For the base case, observe that all nodes within a block of $1$’s form a set of closed twins. Hence, at any second-order stationary point $\boldsymbol{\theta}$, the nodes in the last block $U_{k}$ must synchronize, as follows directly from Corollary 5.6.
Assume as the inductive hypothesis that the nodes in the last $m$ blocks of $1$’s, $U_{k-m+1},\ldots,U_{k}$, are synchronized, i.e., there exists a common phasor $\boldsymbol{v}$ such that $\boldsymbol{v}_{i}=\boldsymbol{v}$ for all $i\in U_{k-m+1}\uplus\cdots\uplus U_{k}$. Lemma 4.4 then implies that the nodes in the adjacent 0-blocks $I_{k-m+1}\uplus\cdots\uplus I_{k}$ also satisfy $\boldsymbol{v}_{i}=\boldsymbol{v}$.
Define